OpenStudy (anonymous):

Please Help!! Find the standard form of the equation of the parabola with a focus at (0, 8) and a directrix at y = -8.

OpenStudy (tkhunny):

The Vertex is exactly between the two given items. Find that, first.

OpenStudy (anonymous):

hmmm....

OpenStudy (anonymous):

u mean write the equation or no @ninjasandtigers

OpenStudy (anonymous):

y = 1/32x^2 y^2 = 8x y^2 = 32x y = 1/8x^2

OpenStudy (anonymous):

between 0 and 8? @tkhunny

OpenStudy (tkhunny):

Well, the problem choices gave it away. Clearly (0,0). Lame.

OpenStudy (tkhunny):

Directrix is y = Something, so this must be vertically oriented. Throw out the ones with y^2.

OpenStudy (tkhunny):

|Directrix to Vertex| = 8 = The Mysterious "p". |Vertex to Focus| = 8 = The Mysterious "p". That's enough information. You can just build it, now.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so is it d?

OpenStudy (aum):

Another way to do this problem is to make use of the property of the parabola that any point on the parabola is equidistant from the focus and the directrix: (x-0)^2 + (y-8)^2 = (y - (-8))^2 x^2 = (y+8)^2 - (y-8)^2 x^2 = (y+8+y-8)(y+8-y+8) x^2 = 2y*16 y = 1/32x^2

OpenStudy (tkhunny):

Always, Always, think about using the definition. It may be more difficult than some specific method, but it ALWAYS works.